OpenStudy (anonymous):

Tina wants to save money for school. Tina invests $1000 in an account that pays an interest rate of 8%. How many years will it take for the account to reach $1300? Round your answer to the nearest hundredth.

OpenStudy (phoenixfire):

Okay, so money is official ridiculous to calculate. Interest rate for this you need a thing called Compound Interest formula which is \[F=P(1+I)^{t}\] F = future amount P = principle amount (initial amount, 1000 in your case) I = interest rate as a decimal (8% = 0.08) t = time in years So you need to solve this equation for t, plug in your numbers, and you'll have your answer. Solving for t: \[1300=1000(1.08)^{t}\] \[{1300 \over 1000} = 1.08^t\] \[\ln 1.3 = t\] \[{{\ln 1.3} \over {\ln 1.08}} = t\] \[t = 3.41\] So your answer is it takes 3.41 years to get $1300 at an interest rate of 8% from an initial amount of $1000